OpenStudy (anonymous):

use properties of exponents to simplify 7 times 7^3 divided by 7^6

OpenStudy (mertsj):

\[\frac{7(7^3)}{7^6}=\frac{7^4}{7^6}=\frac{1}{7^2}\]

OpenStudy (anonymous):

thanks :)

OpenStudy (mertsj):

yw